OpenStudy (anonymous):

1/4 (2x − 14) = 4 x = 15 x = x = 9 x = 11

OpenStudy (callisto):

1/4 (2x − 14) = 4 Multiply both sides by 4 (2x − 14) = 4 x 4 2x - 14 = 16 Add 14 to both sides 2x -14 + 14 = 16+14 2x = 30 Divide both sides by coefficient of x (that is 2 in this case) 2x/2 = 30/2 x = 15